There is a bag filled with 3 blue, 4 red and 5 green marbles.

A marble is taken at random from the bag, the colour is noted and then it is replaced.

Another marble is taken at random.

What is the probability of getting 2 the same colour?



1
Expert's answer
2021-03-10T12:05:39-0500

Let A = {The second marble has the same colour with first taken marble}

"H_i =" {First taken marble has color i} , for "i \in \{red, green, blue\}"

Then according to Law of total probability:

"Pr(A) = \sum_{i \in \{red, green, blue\}}Pr(A | H_i) \cdot Pr(H_i)"

At first, we can state that "Pr(A|H_i) = Pr(H_i)" because marble is replaced, so total count of marbles is equal when we took them first or second time.

We can calculate each of these probablilities using general formula:

"Pr(H_i) =" (count of marbles with color i) / (total count of marbles)

"Pr(H_r) = 4 \/ 12 = 1\/3"

"Pr(H_g) = 5\/12"

"Pr(H_b) = 3\/12 = 1\/4"

"Pr(A) = (1\/3)^2 + (5\/12)^2 + (1\/4)^2 = 50\/144 = 25\/72 = 0.35"

So, answer is 25/72 (or approximately 35%).